# AutoElicit: Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v4

@@ 1 Introduction
Labelling data in medicine can be a costly and time-consuming undertaking (Ghassemi et al., 2020; Cooper et al., 2023). We are particularly interested in the development of automated tools to assist with healthcare monitoring for people living with dementia. In our study, participant data is recorded through passive monitoring devices. Although input features are continuously collected at scale, the process of labelling adverse health conditions, such as urinary tract infection (UTI), can be challenging. Expert knowledge of a task, as privileged information (Vapnik et al., 2015), can provide prior distributions of model parameters, which often reduces the number of samples needed to learn performant models while also enabling more accurate estimates of predictive uncertainty. However, this expert knowledge about model parameters can be difficult to obtain due to the challenges of performing expert prior elicitation, particularly in low-resource settings (Mikkola et al., 2021).
Linear models are widely used in real-world applications, especially in healthcare and as part of our wider study on dementia care, due to their interpretability, robustness, and clinical familiarity (Tennant et al., 2024; Capstick et al., 2024; Ramírez Medina et al., 2025). Many studies do not use prior distributions on their models, possibly due to the challenges associated with elicitation, but would likely benefit greatly from them.
@@
Figure 1: AutoElicit: Prior elicitation using a language model. The left hand figure demonstrates the process, whilst the right hand figure illustrates the benefits of using AutoElicit, achieving the same peak accuracy after $n\approx 15$ labels, 220 days earlier in the study.
In-context learning has been studied as a way to use LLMs for predictive tasks (Dong et al., 2022). Here, LLMs are given a task description and demonstration examples and asked to make predictions on unseen data directly (Brown et al., 2020). There are several reasons why this prompted prediction may not suffice. First, LLMs can incur a large computational cost to run, rendering their use in high-frequency or low-resource settings unlikely. Second, applications may require arithmetic operations within models, an area where LLMs continue to struggle (Spithourakis & Riedel, 2018). Third, LLMs, unless explicitly trained on a specific predictive problem, lack transparency and may not perform Bayesian inference (Min et al., 2022), which can lead to unreliable posterior updates and predictions.
@@
  • We introduce AutoElicit, an algorithm to generate LLM-elicited priors over linear models. We find using LLMs for expert prior elicitation (Figure 1) provide higher posterior accuracy and lower mean square error than an uninformative prior on various numerical predictive tasks, improving models with fewer examples.
  • To compare prior elicitation with in-context learning (ICL), we develop an approach to extract the implicit priors used in ICL which we find to be less informative. To further understand the differences, we study whether an LLM performs Bayesian inference during ICL. We extract the in-context posterior distribution and measure its divergence from posterior Monte Carlo samples of the extracted prior. We find that the language model inconsistently approximates Bayesian inference on real-world datasets.
@@ 2 Related Work
Garthwaite et al. (2005) and O’Hagan et al. (2006) discuss eliciting prior distributions from human experts using a standardised procedure: Experts are trained in probabilities to provide judgements of a task, which can be used to fit probability distributions. Later, Oakley & O’Hagan (2019) provide software and an e-learning course, which trains human experts in the knowledge required to describe informative priors. However, this can be challenging, costly, or infeasible in low-resource settings (Mikkola et al., 2021).
@@
In parallel, research shows that LLMs encode varied domain knowledge, which could be extracted for predictive tasks, possibly in discussions with experts (Gao et al., 2024). Recent works have used LLMs for prior elicitation in multiple ways. In data pre-processing, Choi et al. (2022) and Hollmann et al. (2024) show that LLMs can perform feature selection and engineering respectively, whilst Zhang et al. (2023) and Zhang et al. (2024) use LLMs to specify a modelling pipeline given a task description. This suggests language models might also be able to provide information over the parameters of predictive models directly. In this direction, Gouk & Gao (2024) use LLMs to generate synthetic data from public datasets, which define priors on linear models through a separate likelihood term. However, they do not consider that these datasets might have appeared in the LLM’s training (Appendix A.8), making it challenging to evaluate. Related, Selby et al. (2024) evaluate LLMs for estimating correlations between features and targets, and Zhu & Griffiths (2024) use them to elicit priors for causal learning, proportion estimation, and everyday quantities. However, this knowledge is not used for predictive models. Further related work is given in Appendix A.2.1. In contrast to the literature, we provide a framework for using LLMs to elicit priors over the parameters of linear models directly, which we evaluate on public and private predictive tasks.
@@
LLMs can also be used to produce posterior predictive distributions. In-context learning is an active research area, mainly focused on natural language tasks (Dong et al., 2022), with many works aiming to explain the underlying mechanisms. Relevant to us, Requeima et al. (2024) prompt LLMs with task descriptions and data to generate posterior predictive distributions for regression tasks. Additionally, Hollmann et al. (2022) train a transformer to solve synthetically generated tasks from a prior on tabular datasets and use in-context learning at prediction time. However, these methods sample LLMs for every prediction, which is resource-intensive and potentially unreliable. Additional related work is given in Appendix A.2.2. In our work, we propose a strategy to approximate the in-context prior and posterior distributions when imitating a linear model, which we compare to LLM-elicited priors. Using these, we empirically evaluate whether in-context learning approximates Bayesian inference before using model selection.
@@ 3.1 Prior elicitation with experts
Prior elicitation constructs prior distributions on model parameters, typically through expert judgement (Oakley & O’Hagan, 2019; Gosling, 2018). This can be resource intensive and requires access to one or more domain experts who have been trained in prior elicitation (Mikkola et al., 2021).
@@ 4.1 AutoElicit: Using language models to elicit priors
The posterior predictive distribution on unseen data $(\tilde{x},\tilde{y})$, given an observed dataset $\mathcal{D}$, becomes:
$p_{M,T}(y|\tilde{x},\mathcal{D})\\ =\int_{\Theta}\sum_{k=1}^{K}p_{M,T}(y|\tilde{x},\theta)p_{M,T}(\theta|\mathcal{D},I_{k})p_{M,T}(I_{k})d\theta$ (3)
Where $p_{M,T}(I)$ describes the distribution over task descriptions and $p_{M,T}(\theta|\mathcal{D},I)$ describes the posterior distribution over the parameterisation of a predictive model, which is calculated using the available data $\mathcal{D}$ and the AutoElicit prior over $\theta$.
@@ 4.2 In-context learning
By repeated sampling, we form distributions that estimate $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. These are over the hidden parameters of the in-context model and give us new visibility into the LLM’s predictive process; assuming the model class can be specified in a prompt (Appendix A.4). Using these, we can test whether an LLM is using the elicited prior in-context and if it is approximating Bayesian inference.
@@
4.3 Bayes factor for model selection
@@ 5.1 Experimental setting
We also evaluate AutoElicit on a private dataset not in the LLM’s training, collected as part of our study on Dementia care (Capstick et al., 2024). This contains $10$ features of daily in-home activity and physiology data, and clinically validated labels of positive or negative urinary tract infection (UTI), based on the analysis of a urine sample by a healthcare professional. This dataset contains $78$ people and $371$ positive or negative days of UTI.
@@
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo and GPT-4-turbo, and DeepSeek-R1-32B with 4-bit quantisation. All details are given in Appendix A.7 and other LLMs are tested in Appendix A.15.
@@ 5.2 How informative are AutoElicit priors?
For Heart Disease, uninformative and GPT-3.5-turbo-elicited priors produced the same mean posterior accuracy, as this elicited prior contains many standard Normal distributions (Figure 15). This is reassuring, as GPT-3.5-turbo responded with an uninformative prior when unsure of its knowledge. However, GPT-4-turbo and DeepSeek-R1-32B-Q4, returned considerably better priors. The appropriate prior can be chosen with Bayes factor (Appendix A.16)
@@
These results illustrate the remarkable effectiveness of using AutoElicit priors, which considerably improve posterior predictions, particularly for smaller training sizes; where label collection might have been challenging. Our framework makes expert prior elicitation more accessible.
@@ 5.3 Can experts supplement the process?
Including expert information in the task description allows AutoElicit priors to combine learnt insights with domain knowledge (Capstick et al., 2024). In the UTI task description we now list three features that are positively correlated with the label: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
Figure 3 shows that this information is incorporated when LLMs provide priors, as these three features become more positively predictive of a UTI. GPT-3.5-turbo had already identified this relationship, shown by the positive median of the priors, which explains the small improvement in posterior accuracy this additional information provides. Further, Appendix A.11, studies the priors elicited as we incrementally improve the description of a synthetic task, illustrating that GPT-3.5-turbo produces reliable distributions from text.
Figure 3: UTI prior parameters with expert information. The histograms in the three left hand plots show the distribution of parameter values for the features that we provide expert information for in the task description. In all three cases, we state that the features are positively correlated with UTI risk. In the right hand plot, we show the posterior accuracy of these distributions for different numbers of observed samples. The average and 95% confidence interval of the mean posterior accuracy over the $10$ test splits are shown.
@@ 5.4 The in-context priors and posteriors
Figures 4 and 5 show GPT-3.5-turbo’s in-context prior and posterior distributions approximated with maximum likelihood estimation (MLE). To quantify the error, we calculate the mean squared error (MSE) between the LLM’s in-context predictions and the approximated model’s predictions, shown in the bottom row of Figures 4 and 5. This allows us to study the implicit parameters used by the LLM to make predictions in-context.
@@
The estimated in-context posteriors in Figure 5 differ from the priors, suggesting the LLM has updated its knowledge based on demonstrations. For the Diabetes and Heart Disease tasks, the LLM mostly continued to produce linear predictions, and so our error increases, but remains low. However, the Breast Cancer and Hypothyroid dataset produce a large approximation error, as the LLM struggled to imitate a linear model. The approximation error on the synthetic task is discussed in Appendix A.4. We do not estimate the in-context posterior for the UTI dataset as it is private, and so we are unable to provide demonstrations.
@@ 5.5 Is the language model consistent?
Table 1: Approximated GPT-3.5-turbo in-context prior compared to the elicited prior. The energy statistic between the LLM’s approximate in-context prior, and the LLM-elicited prior. Lower values indicate that the two distributions are more similar.
@@ 5.6 Does in-context learning use Bayesian inference?
These two posterior samples are shown in Figure 6, where we focus on those datasets where in-context learning produced linear predictions, providing a low posterior approximation error. For the Diabetes task, samples from the approximated in-context posterior and MC sampling on the approximated in-context prior visually agree and produce a low energy statistic, suggesting the LLM is performing Bayesian inference in-context. However, for the Heart Disease task the two posterior samples differ, showing that the LLM inconsistently applies the Bayes rule in-context. Results of the other datasets and training splits are given in Appendix A.14. This suggests LLM-elicited priors are more reliable, particularly when access to the posterior distribution is important, such as for uncertainty estimation.
Figure 6: The in-context posterior distribution. We compare the posterior extracted from GPT-3.5-turbo’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. For the diabetes dataset we have removed the long tails for improved visualisation.
@@
5.7 Using Bayes factor for model selection
If in-context learning approximates Bayesian inference, we calculate the Bayes factor between it and AutoElicit priors.
Figure 7: The distribution of the prior marginal likelihood and the Bayes factor for GPT-3.5-turbo. This shows the distribution of prior likelihoods, calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is calculated between prior elicitation and in-context learning, with a number greater than $1$ favouring prior elicitation. For the regression tasks, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles in the histograms.
Figure 7 shows the marginal likelihood distribution for in-context learning, AutoElicit, and an uninformative prior along with the Bayes factor (where reliably defined). Unlike in-context learning, AutoElicit priors are the better choice over an uninformative prior for all datasets tested. Whether AutoElicit or in-context learning provides better priors depends on the task. However, in only two of the five datasets tested, in-context learning reliably produced linear predictions, and in only one of those it approximated Bayesian inference. Considering this, the cost, and the ease of interpretability, AutoElicit is likely the better option.
This model selection strategy is designed to be more relevant as LLMs become more reliable, and their cost reduces.
@@ 6 Limitations and future research
Finally, it would be interesting to extend our results to neural networks. An immediate option is to build on the literature asking the LLM to suggest data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Further, extending our results to non-linear models such as random forests or Gaussian processes would be valuable.

### v5

@@ Abstract
LLMs for healthcare, Prior Elicitation, Large Language Models
[…]
sas
@@ 1 Introduction
Labelling data in medicine can be a costly and time-consuming undertaking (Ghassemi et al., 2020; Cooper et al., 2023). We are particularly interested in the development of automated tools to assist with healthcare monitoring for people living with dementia. In our study, participant data is recorded through passive monitoring devices. Although input features are continuously collected at scale, the process of labelling adverse health conditions, such as urinary tract infection (UTI), can be challenging. Expert knowledge of a task, as privileged information (Vapnik et al., 2015), can provide prior distributions of model parameters, which often reduces the number of samples needed to learn performant models while also enabling more accurate estimates of predictive uncertainty. However, this expert knowledge about model parameters can be difficult to obtain due to the challenges of performing expert prior elicitation, particularly in low-resource settings (Mikkola et al., 2024).
Linear models are widely used in real-world applications, especially in healthcare and as part of our wider study on dementia care, due to their interpretability, robustness, and clinical familiarity (Capstick et al., 2024; Tennant et al., 2024; Ramírez Medina et al., 2025). Many studies do not use prior distributions on their models, possibly due to the challenges associated with elicitation, but would likely benefit greatly from them.
@@
Figure 1: AutoElicit: Prior elicitation using a language model. The left-hand figure demonstrates the process, whilst the right-hand figure illustrates the benefits of using AutoElicit, achieving the same peak accuracy after $n\approx 15$ labels, 220 days earlier in the study.
In-context learning has been studied as a way to use LLMs for predictive tasks (Dong et al., 2024). Here, LLMs are given a task description and demonstration examples and asked to make predictions on unseen data directly (Brown et al., 2020). There are several reasons why this prompted prediction may not suffice. First, LLMs can incur a large computational cost to run, rendering their use in high-frequency or low-resource settings unlikely. Second, applications may require arithmetic operations within models, an area where LLMs continue to struggle (Spithourakis & Riedel, 2018). Third, LLMs, unless explicitly trained on a specific predictive problem, lack transparency and may not perform Bayesian inference (Min et al., 2022), which can lead to unreliable posterior updates and predictions.
@@
  • We introduce AutoElicit, an algorithm to generate LLM-elicited priors for linear models. We find using LLMs for expert prior elicitation (Figure 1) provides higher posterior accuracy and lower mean squared error than uninformative priors on various numerical predictive tasks, improving models with fewer examples.
  • To compare prior elicitation with in-context learning (ICL), we develop an approach to extract the implicit priors used in ICL, which we find to be less informative. To further understand the differences, we study whether an LLM performs Bayesian inference during ICL. We extract the in-context posterior distribution and measure its divergence from posterior Monte Carlo samples of the extracted prior. We find that the language model inconsistently approximates Bayesian inference on real-world datasets.
@@ 2 Related Work
Garthwaite et al. (2005) and O’Hagan et al. (2006) discuss eliciting prior distributions from human experts using a standardised procedure: Experts are trained in probabilities to provide judgements of a task, which can be used to fit probability distributions. Later, Oakley & O’Hagan (2019) provide software and an e-learning course, which trains human experts in the knowledge required to describe informative priors. However, this can be challenging, costly, or infeasible in low-resource settings (Mikkola et al., 2024).
@@
In parallel, research shows that LLMs encode varied domain knowledge, which could be extracted for predictive tasks, possibly in discussions with experts (Gao et al., 2024). Recent works have used LLMs for prior elicitation in multiple ways. In data pre-processing, Choi et al. (2022) and Hollmann et al. (2024) show that LLMs can perform feature selection and engineering respectively, whilst Zhang et al. (2023) and Zhang et al. (2024) use LLMs to specify a modelling pipeline given a task description. This suggests language models might also be able to provide information over the parameters of predictive models directly. In this direction, Gouk & Gao (2024) use LLMs to generate synthetic data from public datasets, which define priors on linear models through a separate likelihood term. However, they do not consider that these datasets might have appeared in the LLM’s training (Appendix A.8 and A.9.5), making it challenging to evaluate. Related, Selby et al. (2025) evaluate LLMs for estimating correlations between features and targets, and Zhu & Griffiths (2024) use them to elicit priors for causal learning, proportion estimation, and everyday quantities. However, this knowledge is not used for predictive models. Further related work is discussed in Appendix A.2.1. In contrast, we provide a framework for using LLMs to elicit priors over the parameters of linear models directly, which we evaluate on public and private predictive tasks.
@@
LLMs can also be used to produce posterior predictive distributions. In-context learning is an active research area, mainly focused on natural language tasks (Dong et al., 2024), with many works aiming to explain the underlying mechanisms. Relevant to us, Requeima et al. (2024) prompt LLMs with task descriptions and data to generate posterior predictive distributions for regression tasks. Additionally, Hollmann et al. (2022, 2025) train a transformer to solve synthetically generated tasks from a prior on tabular datasets and use in-context learning at prediction time. However, these methods sample LLMs for every prediction, which is resource-intensive and potentially unreliable. Additional related work is given in Appendix A.2.2. In our work, we propose a strategy to approximate the in-context prior and posterior distributions when imitating a linear model, which we compare to LLM-elicited priors. Using these, we empirically evaluate whether in-context learning approximates Bayesian inference before using model selection.
@@ 3.1 Prior Elicitation with Experts
Prior elicitation constructs prior distributions on model parameters, typically through expert judgement (Oakley & O’Hagan, 2019; Gosling, 2018). This can be resource-intensive and requires access to one or more domain experts who have been trained in prior elicitation (Mikkola et al., 2024).
@@ 4.2 In-context Learning
By repeated sampling, we form distributions that estimate $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. These are over the hidden parameters of the in-context model and give us new visibility into the LLM’s predictive process, assuming the model class can be specified in a prompt (Appendix A.4). Using these, we can test whether an LLM is using the elicited prior in-context and if it is approximating Bayesian inference.
@@
4.3 The Bayes Factor for Model Selection
@@ 5.1 Experimental Setting
We also evaluate AutoElicit on a private dataset not in the LLM’s training, collected as part of our study on Dementia care. This contains $10$ features of daily in-home activity and physiology data, and clinically validated labels of positive or negative urinary tract infection (UTI), based on the analysis of a urine sample by a healthcare professional. This dataset contains $78$ people and $371$ positive or negative days of UTI.
@@
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo and GPT-4-turbo, and DeepSeek-R1-32B with 4-bit quantisation. All details are given in Appendix A.7, and other LLMs are tested in Appendix A.16.
@@ 5.2 How Informative are AutoElicit Priors?
For Heart Disease, uninformative and GPT-3.5-turbo-elicited priors produced the same mean posterior accuracy, as this elicited prior contains many standard Normal distributions (Figure 18). This is reassuring, as GPT-3.5-turbo responded with an uninformative prior when unsure of its knowledge. However, GPT-4-turbo and DeepSeek-R1-32B-Q4, returned considerably better priors. The appropriate prior can be chosen with the Bayes factor (Appendix A.17).
@@
In Appendix A.9.5, we also compare our results to those achieved using the approach presented by Gouk & Gao (2024). This method uses an LLM to generate feature values before labelling them with zero-shot in-context learning. These generated samples are provided alongside real data during the training of a linear predictive model. Therefore, the method presented in Gouk & Gao (2024) is susceptible to LLM data memorisation, putting it at risk of artificially providing improved posterior results on public datasets. In light of this, Figure 13 shows the results split by those on private and public tasks. Our approach yields improved posterior accuracy and mean squared error over this baseline.
In Appendix A.13, we demonstrate how the priors elicited can be useful for explaining the final predictive model.
These results illustrate the remarkable effectiveness of using AutoElicit priors, which considerably improve posterior predictions, particularly for smaller training sizes, where label collection might have been challenging. Our framework makes expert prior elicitation more accessible.
@@ 5.3 Can Experts Supplement the Process?
Including expert information in the task description allows AutoElicit priors to combine learnt insights with domain knowledge. In the UTI task description, we now list three features that positively correlate with the label (Capstick et al., 2024): “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
Figure 3 shows that this information is incorporated when LLMs provide priors, as these three features become more positively predictive of a UTI. GPT-3.5-turbo had already identified this relationship, shown by the positive median of the priors, which explains the small improvement in posterior accuracy this additional information provides. Further, Appendix A.11 studies the priors elicited as we incrementally improve the description of a synthetic task, illustrating that GPT-3.5-turbo produces reliable distributions from text.
Figure 3: UTI prior parameters with expert information. The histograms in the three left-hand plots show the distribution of parameter values for the features that we provide expert information for in the task description. In all three cases, we state that the features are positively correlated with UTI risk. In the right-hand plot, we show the posterior accuracy of these distributions for different numbers of observed samples. The average and 95% confidence interval of the mean posterior accuracy over the $10$ test splits are shown.
@@ 5.4 The In-context Priors and Posteriors
Figures 4 and 5 show GPT-3.5-turbo’s in-context prior and posterior distributions approximated with maximum likelihood estimation (MLE). This allows us to study the implicit parameters used by the LLM to make predictions in-context.
@@
The estimated in-context posteriors in Figure 5 differ from the priors, suggesting the LLM has updated its knowledge based on demonstrations. For the Diabetes and Heart Disease tasks, the LLM mostly continued to produce linear predictions, and so our error increases, but remains low. However, the Breast Cancer and Hypothyroid dataset produces a large approximation error, as the LLM struggles to imitate a linear model. The approximation error on the synthetic task is discussed in Appendix A.4. We do not estimate the in-context posterior for the UTI dataset as it is private, and so we are unable to provide demonstrations.
@@ 5.5 Is the Language Model Consistent?
Table 1: Approximated GPT-3.5-turbo in-context prior compared to the elicited prior. The energy statistic between the LLM’s approximate in-context prior and the LLM-elicited prior. Lower values indicate that the two distributions are more similar.
@@ 5.6 Does In-context Learning use Bayesian Inference?
These two posterior samples are shown in Figure 6, where we focus on those datasets where in-context learning produced linear predictions, providing a low posterior approximation error. For the Diabetes task, samples from the approximated in-context posterior and MC sampling on the approximated in-context prior visually agree and produce a low energy statistic, suggesting the LLM is performing Bayesian inference in-context. However, for the Heart Disease task, the two posterior samples differ, showing that the LLM inconsistently applies the Bayes rule in-context. Results of the other datasets and training splits are given in Appendix A.15. This suggests LLM-elicited priors are more reliable, particularly when access to the posterior distribution is important, such as for uncertainty estimation.
Figure 6: The in-context posterior distribution. We compare the posterior extracted from GPT-3.5-turbo’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. For the diabetes dataset, we have removed the long tails for improved visualisation.
@@
5.7 Using the Bayes Factor for Model Selection
If in-context learning approximates Bayesian inference, we can calculate the Bayes factor between it and AutoElicit.
Figure 7: The distribution of the prior marginal likelihood and the Bayes factor for GPT-3.5-turbo. This shows the distribution of prior likelihoods, calculated using $25$ data points from each dataset, over $5$ test splits and $100$ samples of prompts. The Bayes factor is calculated between prior elicitation and in-context learning, with a number greater than $1$ favouring prior elicitation. For the regression tasks, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles in the histograms and assume the priors form a single Gaussian.
Figure 7 shows the marginal likelihood distribution for in-context learning, AutoElicit, and an uninformative prior, along with the Bayes factor (where reliably defined). Unlike in-context learning, AutoElicit priors are the better choice over an uninformative prior for all datasets tested. Whether AutoElicit or in-context learning provides better priors depends on the task. However, in only two of the five datasets tested, in-context learning reliably produced linear predictions, and in only one of those it approximated Bayesian inference. Considering this, the cost, and the ease of interpretability, AutoElicit is likely the better option.
This model selection strategy is designed to be more relevant as LLMs become more reliable and their costs reduce.
@@ 6 Limitations and Future Research
Finally, it would be interesting to extend our results to neural networks. An immediate option is to build on the literature, asking the LLM to suggest data augmentation techniques for a task, effectively reducing the cost of data collection by extending the current dataset. Further, extending our results to non-linear models such as random forests or Gaussian processes would be valuable.